OpenStudy (anonymous):

which of the following accounts would have the largest total return after 1 year? a) 6% APR compounded monthly b) 6% APR compounded annually c) 6% APY d) 6% APR compounded continuously e) not enough information to tell f) they all would have the same total return

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

for the formula for compounded continuously is A = Pe^rt where P is your principal and r is your annual rate and t is year. monthly and annually can be calculated using A = P (1+ r/n)^nt where n is the times compounded annually and r is interest rate (in decimal) and t is time in years.. if you plug in some dummy value for P you can check each one.

OpenStudy (anonymous):

example \[A = 100*e ^{0.06*1}\] which should give you the return on 100$ invested for 1 year at 0.06 (6%) continuously for 1 year